Compute the number of functions f from the set {0,1,2,...; n} (where n is a positive integer), to the set {0,1}.
1
Expert's answer
2020-11-11T19:23:45-0500

Let us compute the number of functions "f:\{0,1,...,n\}\to\{0,1\}". For each "k\in\{0,1,...,n\}" for the value "f(k)" there are two possibility: "f(k)=0" or "f(k)=1". Since the cardinality of the set "\{0,1,...,n\}" is "n+1", by Multiplication Principle the number of all functions is

"\underbrace{2\cdot 2\cdot ...\cdot 2}_{n+1}= 2^{n+1}."